A jet traveled 1029 km in 6 hours flying aginst a headwind. The return trip (with the headwind) took 3 1/3 hours. Find the speed of the jet in still air.
Answer by checkley75(3666)

6R=1029
R=1029/6
R=171.5 SPEED AGAINST THE HEADWIND
1029/3.333=R
R=308.73 IS THE SPEED WITH THE HEADWIND
THE DIFFERENCE BETWEEN THESE TWO SPEEDS DIVIDED BY 2 IS THE HEADWIND
(308.73-171.5)/2=137.23/2=68.62 IS THE SPEED OF THE HEADWIND.